OpenStudy (anonymous):

suppose that y varies inversely with x, and y = 6 when x = 8. what is an equation for the inverse variation?

OpenStudy (skullpatrol):

Any ideas?

Parth (parthkohli):

Remember that the equation for inverse variation is \(k = xy\).

OpenStudy (aravindg):

\[\large y \alpha \dfrac{1}{x}\]

Parth (parthkohli):

Or, equivalently, it can be expressed as\[y = \dfrac{k}{x}\]

Parth (parthkohli):

@AravindG use `\propto` instead of `\alpha` ;-)

OpenStudy (skullpatrol):

Do you know these general forms of an inverse relation?

Parth (parthkohli):

\[k = yx\]Here, we can plug the values \(y = 6\) and \(x = 8\). What do you think?

OpenStudy (anonymous):

woah o.o i was doing something else. K=6(8)

Parth (parthkohli):

I'd give you an example. There's an inverse relation and when \(y = 4\), \(x = 6\). The general equation would be \(k = yx\). find out \(k\), the product of all \(y\)s and the corresponding \(x\)'s.

Parth (parthkohli):

So the general equation for my example is \(k = 24\)

Parth (parthkohli):

Use my example?

OpenStudy (anonymous):

i wish i had time to learnt his but i dont. 20 questions to go and i have to turn this in at @ 2. its 1:22 now -.-

Parth (parthkohli):

It's very easy. What's the product of the \(x\) and \(y\) you're given?

OpenStudy (anonymous):

y = 48/x ?

Parth (parthkohli):

Yay! That's it! :-)

OpenStudy (anonymous):

ಥ◡ಥ